• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 08:03:02 on September 19, 2005

    Entry number 155177

    keyword=Owl Shift Summary

    Owl Shift for Monday September 19th, 2005


    Leader:   Bryan Moffit
    Worker/Target:   Vince Sulkosky

    Left HRS p0   = -2.737 GeV/c   Septum Set Current = 283.57A
    Right HRS p0  = -2.733 GeV/c   Septum Set Current = 278.22A
    Sweeper is ON at 111.00A

    Summary

    All production this shift (production runs 5169-5177: 383.3k pairs ~= 1.02C).

    Significant Events

    • 00:00 Began shift with the start of run 5168, compton run 12222. IHWP IN.
    • 00:08 Asked MCC to increase beam current to 40µA (from 39µA).
    • 00:15 Vince and I notice that the Hall-C charge asymmetry is huge. After talking with MCC and reading the ELOG, determined that Hall-C charge feedback will not work while they are using the Hall-B laser. Bummer.
    • 03:12 Asked MCC to retract the IHWP.
    • 07:54 Asked MCC to insert the IHWP.

    • Data Run Summary

      • 23:18 Run 5168 - Production Run (20cm Helium-4), IHWP IN, 42.4k pairs.
      • 00:15 Run 5169 - Production Run (20cm Helium-4), IHWP IN, 5.3k pairs.
      • 00:22 Run 5170 - Production Run (20cm Helium-4), IHWP IN, 46.8k pairs.
      • 01:19 Run 5171 - Production Run (20cm Helium-4), IHWP IN, 49.0k pairs.
      • 02:15 Run 5172 - Production Run (20cm Helium-4), IHWP IN, 46.9k pairs.
      • 03:13 Run 5173 - Production Run (20cm Helium-4), IHWP OUT, 47.8k pairs.
      • 04:09 Run 5174 - Production Run (20cm Helium-4), IHWP OUT, 48.0k pairs.
      • 05:06 Run 5175 - Production Run (20cm Helium-4), IHWP OUT, 48.9k pairs.
      • 06:02 Run 5176 - Production Run (20cm Helium-4), IHWP OUT, 46.7k pairs.
      • 06:58 Run 5177 - Production Run (20cm Helium-4), IHWP OUT, 43.9k pairs.
      • 07:55 Run 5178 - Production Run (20cm Helium-4), IHWP IN. Passed to day shift.

      Compton Data Run Summary

      • 00:00 Run 12222 - IHWP IN. Beam polarization = -84.1%,
      • 00:09 Run 12223 - IHWP IN, ROC5 crashed.
      • 02:19 Run 12224 - IHWP IN, Beam polarization = -85.7%,
      • 03:15 Run 12225 - IHWP OUT, Beam polarization = 83.7%, ROC5 Crashed.
      • 05:12 Run 12226 - IHWP OUT, Junk. No rates.
      • 05:13 Run 12227 - IHWP OUT, Junk. No rates.
      • 05:15 Run 12228 - IHWP OUT, Beam polarization = 83.3%,
      • 07:14 Run 12229 - IHWP OUT, Beam polarization = ??.?%, ROC1 crashed.
      • 07:38 Run 12230 - IHWP OUT. No coincidences.
      • 07:55 Run 12231 - IHWP IN. Passed to day shift.